OpenStudy (agentnao):

An oil tanker spills 100,000 cubic meters of oil, which forms a lick that spreads on the water surface in a shape best modeled by a circular disc of radius r meters and thickness h meters. The radius of the circular disc is increasing at a rate of 3m/min. At t=T, the area of the "circular" slick reaches 100 pi m^2. a) how fast is the area of the slick increasing at t=T? b) how fast is the thickness of the slick decreasing at t=T? c) Find the rate of change of the area of the slick with respect to the thickness at t=T?

OpenStudy (anonymous):

(a) Now, A = pi*r^2 dA/dt = 2*pi*r*dr/dt so at time t, when area is 100 * pi m^2 , radius = 10m and dr/dt = 3 (thats given) so, dA/dt = 2 * pi * 10 * 3 m^2 / min therefore, dA/dt = 60 pi m^2/ min (b) Disk Volume = Area * Thickness Volume is constant (assuming density of oil won't change and mass of oil spill was fixed) so d/dt(Volume) = 0 = hdA/dt + Adh/dt therefore, dh/dt = -h/A * dA/dt Now at time = t, area = 100 * pi m^2 therefore h = 100000/(100 * pi) = 1000/pi m hence dh/dt = - 1000/pi * 1/(100 * pi) * 60 * pi = - 600/pi m / min hence rate of decrease of h = 600/pi m /min (c) Disk Volume = Area * Thickness Volume is constant (assuming density of oil won't change and mass of oil spill was fixed) so d/dt(Volume) = 0 = hdA/dt + Adh/dt therefore, dh/dt = -h/A * dA/dt

OpenStudy (agentnao):

I don't really understand part c, but I get what you did for the other two! Thank you so much! :D

OpenStudy (anonymous):

you are right, for part c , we have to find dA/dh . so V = A * h differentiate with respect to h dV/dh = hdA/dh + A and dV/dh = 0 (because Volume is essentially constant) hence hdA/dh + A = 0 dA/dh = -A/h

OpenStudy (perl):

I think for part c) you can use the fact that dA/dh = ( dA/dt ) / ( dh/dt)

OpenStudy (perl):

@zaphodplaysitsafe are you sure about your answer. I get dA/dh = (60 pi ) / ( -600 /pi ) = - 1/10 * pi^2